OpenStudy (anonymous):

Solve the following quadratic equation: -2x^2 – x – 1 = 0

OpenStudy (anonymous):

\[ 2x^2 + x –1 = 0\\ a=2; b=1; c=1;\\ b^2 - 4ac= 1 -8=-7\\ \sqrt{-7}= i\sqrt 7 \]

OpenStudy (anonymous):

\[ x= \frac{1}{4} \left(-1\pm i \sqrt{7}\right) \]